# Existence proofs of traveling wave solutions on an infinite strip for the suspension bridge equation and proof of orbital stability

### v1

@@ 5 Stability Analysis
In this section, we aim at investigating the stability of traveling wave solutions to (2), of which the existence will be constructively established in Section 6. In particular, we assume that we know the existence of a solution $\tilde{u}\in\mathcal{H}$ to (2) such that
@@
$E\begin{pmatrix}u\\ v\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}\frac{1}{2}v(x)^{2}+\frac{1}{2}(\Delta u(x))^{2}+e^{u(x)}-u(x)-1dx$ (37)
@@
Let $\tilde{u}$ be a traveling wave solution to (2) in $\mathcal{H}$ and let $\tilde{U}$ be defined as in (39). Then $\tilde{u}$ is said to be orbitally stable if the solution $U$ to (38) with initial condition $U(0)=U_{0}$ exists for all positive times and if for all $\epsilon>0$, there exists $\delta>0$ such that if $\|U_{0}-\tilde{U}\|_{X}<\delta$, then $\inf_{s\in{\mathbb{R}}}\|U(t)-T(s)\tilde{U}\|_{X}<\epsilon.$
$\tilde{u}$ is called orbitally unstable if it is not orbitally stable.
@@
Using the notations of [18], $\tilde{w}$ in Lemma 5.1 corresponds to $v_{c}$ in Section 4.3 of [18]. Moreover, $\theta$ corresponds to $d^{\prime\prime}(c)$, after using integrations by parts (cf. Section 4.3 of [18]).
@@
This concludes the proof. ∎
@@
for some invertible infinite matrix $P=P^{N}+\pi_{N}$, with $P^{N}=\pi^{N}P^{N}\pi^{N}$, whose inverse is given as $P^{-1}=\pi^{N}(P^{N})^{-1}\pi^{N}+\pi_{N}$. $(P^{N})^{-1}$ has to be understood as the inverse of $P^{N}:\pi^{N}\ell^{2}\to\pi^{N}\ell^{2}$. In practice, columns of $P$ are approximate eigenvectors of $DF(\bar{U})$, so that $\mathcal{D}$ is close to being diagonal. We define $S$ as the diagonal part of the matrix $\mathcal{D}$ and $R=\mathcal{D}-S$. Moreover, we define $\lambda_{k}$ as the diagonal entries of $S$, that is
@@
Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}(\mathcal{J})$, $b_{n_{2}}(\mathcal{J})$ and ${C}_{n_{2}}(\mathcal{J})$ as
@@
Moreover, given $\mu\in\mathcal{J}$, we define $f_{n_{2},\mu}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}(\frac{1}{l_{n_{2}}-\mu})$. Then, we obtain that
@@
Then, one can prove that both $a_{n_{2}}$ and $b_{n_{2}}$ are decreasing functions of $\mu$, leading to the desired result. ∎
@@
Let $t>-\lambda_{min}$ and let $\mathcal{J}=[\lambda_{min},\delta_{0}]$. Moreover, let us define $E=(E_{n})_{n\in\mathbb{N}_{0}^{2}}$ as follows:
@@
Now, let us introduce various constants satisfying the following
@@
we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ If in addition
$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0,$ (53)
then we define
@@
The proof of the lemma follows the ones of Lemmas 4.1 and 4.2 in [4]. Now, we comment on the computation of the bounds. Following Lemma 4.1 in [4], we want to compute $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ such that
@@
Now, noticing that $\left(\mathbb{L}_{n_{2}}-\mu I\right)^{-1}$ is the convolution operator associated to the function $f_{n_{2},\mu}$ defined in Lemma 5.4, we combine Lemma 4.7 and Lemma 5.4 to obtain that
$2d_{1}(\bar{V},\bar{V}*E)$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2}$
@@
using Lemma 7.1. Then, recalling $C_{n_{2}}(\mu)$ from the proof of Lemma 5.4, one can prove that $\mu\mapsto C_{n_{2}}(\mu)\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$ is increasing in $\mu$ and consequently the maximum is reached at $\mu=\delta_{0}$ when considering $\mu\in\mathcal{J}$. This concludes the obtained inequality for $\mathcal{Z}_{u,3}$. Now, the computations of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ follow directly from the proof of Lemma 4.3. In particular, we also use that $\|(\mathbb{L}-\mu I)^{-1}\|_{2}\leq\frac{1}{1-\frac{c^{4}}{4}-\delta_{0}}$ for all $\mu\in\mathcal{J}$. The rest of the proof follows from [4]. ∎
@@ 5.2 Enclosure of $\theta$
Let $r_{0}>0$ and let $\tilde{u}\in\mathcal{H}$ be a solution to (2) such that $\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}$, where $\bar{u}\in\mathcal{H}$ is constructed as in (20). Let $\overline{W}\in\pi^{N}\ell^{2}$ such that $\overline{w}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\overline{W})\in\mathcal{H}$. Let $\epsilon_{0},\epsilon>0$ be defined as
@@
where $\overline{w}\in\mathcal{H}$ is an approximation of $\tilde{w}$. In particular, using Cauchy-Schwarz inequality, we have
@@
using that $\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}=\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}.$ Moreover, using the Fourier transform, we have that
@@
On the other hand, we have
@@
$\leq(\kappa_{1}r_{0}+2c\|\tilde{w}-\overline{w}\|_{2})(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\frac{r_{0}}{2-c^{2}}).$
Consequently, it remains to compute $\|\tilde{w}-\overline{w}\|_{2}$. By definition of $\tilde{w}$ in (40), we have
$\|\tilde{w}-\overline{w}\|_{2}=\|-2cD\mathbb{F}(\tilde{u})\partial_{x_{1}}^{2}\tilde{u}-\overline{w}\|_{2}\leq 2c\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
Now, we have
$\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}\leq\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\|\partial_{x_{1}}^{2}(\bar{u}-\tilde{u})\|_{2}+\frac{1}{2c}\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
Moreover, we obtain that
@@
Finally, using that $\bar{u}=\gamma^{\dagger}(\bar{U})$ and $\overline{w}=\gamma^{\dagger}(\overline{W})$, we use Parseval’s identity to get
@@
where $(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}$ for all $(n_{1},n_{2})\in\mathbb{N}_{0}^{2}$. This concludes the proof. ∎
@@ 6 Results
In this section, we present the main results of the paper, which consist of rigorous, computer-assisted bounds on numerical approximations to (2). These approximate solutions were obtained using Newton’s method, initialized with carefully constructed guesses based on approximations in the periodic setting described in [32]. To extend the initial guesses in the periodic case to the infinite setting, we project them into the kernel of matrix $\mathcal{T}$, as described in Section 3.1.
@@
Recall that $\nu$ and $N^{\mathrm{FFT}}$ are defined in Section 3.2. We choose $\nu=(1.1,1.1)$ and $N^{\mathrm{FFT}}=(512,512)$. Finitely many cosine coefficients of the approximate solution $\bar{u}$, visualized in Figure 1, can be found in the code in [31]. We report the values for the bounds in floating point numbers displaying four decimals. We have
@@
The stability part follows similarly as the one presented in the proof of Theorem 6.1. In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from on another, and disjoint from the rest of the intervals. ∎
@@
For the stability part, we obtain $\lambda_{min}\approx-15$, using the notation of Lemma 5.3. Then, we choose $t>\lambda_{min}$ in Lemma 5.5 and compute the bounds of the lemma to compute the Gershgorin intervals $\overline{B_{\epsilon_{j}}(\lambda_{j})}$. We obtained intersecting intervals and could not conclude about the localization of negative eigenvalues. However, we obtained that the union of all the intervals was contained in $(-0.4,\infty)$. This implies that we can improve our choice for $\lambda_{min}$ and choose $\lambda_{min}=-0.45$. Choosing $t=0.5$ in Lemma 5.3 and computing the bounds again, we get tighter radiae for the Gershgoring intervals and obtain disjoint intervals on the negative part of $\mathbb{R}$}.
@@ 7 Appendix
In the next lemma we provide some technical results about $l_{n}$.
@@
$2\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{(2d_{2})^{7}}{6N^{6}}\frac{5}{8(2\pi)^{7}}$
where $N=\lceil{\frac{d_{2}c}{\pi}}\rceil.$

### v3

@@ 5 Stability Analysis
In this section, we aim at investigating the stability of traveling wave solutions to (2) with Neumann boundary conditions (3), which is the traveling wave formulation associated with (1). The existence of these traveling waves will be constructively established in Section 6. In particular, we assume that we know the existence of a solution $\tilde{u}\in\mathcal{H}$ to (2) such that
@@
$E\begin{pmatrix}u\\ v\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}\left[\frac{1}{2}v(x)^{2}+\frac{1}{2}(\Delta u(x))^{2}+e^{u(x)}-u(x)-1\right]dx$ (37)
@@
Let $\tilde{u}$ be a traveling wave solution to (2) in $\mathcal{H}$ and let $\tilde{U}$ be defined as in (39). Then $\tilde{u}$ is said to be orbitally stable if the solution $U$ to (38) with initial condition $U(0)=U_{0}$ exists for all positive times and if for all $\epsilon>0$, there exists $\delta>0$ such that if $\|U_{0}-\tilde{U}\|_{X}<\delta$, then $\inf_{s\in{\mathbb{R}}}\|U(t)-T(s)\tilde{U}\|_{X}<\epsilon.$ Furthermore, $\tilde{u}$ is called orbitally unstable if it is not orbitally stable.
@@
A solution is not only valid for a fixed parameter $c$, but also persists under variations of $c$. Differentiating with respect to $c$ yields $\tilde{w}$ in Lemma 5.1, which is called $v_{c}$ in [18]. To determine (in)stability, the authors of [18] introduce a functional $d(c)$ and compute its second derivative $d^{\prime\prime}(c)$ in Section 4.3 of their paper, which corresponds to $\theta$ in our Lemma 5.1 after applying integrations by parts.
@@
This concludes the proof after a short algebraic manipulation. ∎
@@
for some invertible infinite matrix $P=P^{N}+\pi_{N}$, with $P^{N}=\pi^{N}P^{N}\pi^{N}$, whose inverse is given as $P^{-1}=\pi^{N}(P^{N})^{-1}\pi^{N}+\pi_{N}$. $(P^{N})^{-1}$ has to be understood as the inverse of $P^{N}:\pi^{N}\ell^{2}\to\pi^{N}\ell^{2}$. In practice, columns of $P$ are approximate eigenvectors of $DF(\bar{U})$, so that $\mathcal{D}$ is close to being diagonal. We define $S$ as the diagonal part of the matrix $\mathcal{D}$ and $R=\mathcal{D}-S$. Moreover, we define $\lambda_{k}$ as the diagonal entries of $S$, that is,
@@
Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}(\mathcal{J})$, $b_{n_{2}}(\mathcal{J})$ and ${C}_{n_{2}}(\mathcal{J})$ as (cf. Lemma 4.6)
@@
Given $\mu\in\mathcal{J}$, we define $f_{n_{2},\mu}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}(\frac{1}{l_{n_{2}}-\mu})$. Then, we obtain that
@@
Subsequently, one can prove that both $a_{n_{2}}$ and $b_{n_{2}}$ are decreasing functions of $\mu$, leading to the desired result. ∎
@@
Let $t>-\lambda_{min}$ and let $\mathcal{J}=[\lambda_{min},\delta_{0}]$. We define $E=(E_{n})_{n\in\mathbb{N}_{0}^{2}}$ as
@@
We introduce various constants satisfying the following inequalities:
@@
we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ Furthermore, we assume that
$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0.$ (53)
Then, we define
@@
for all $n\in\mathbb{N}_{0}^{2}$. Let $k\in\mathbb{N}$ and $I\subset\mathbb{N}_{0}^{2}$ such that $|I|=k$.
@@
The proof of the lemma follows the ones of Lemmas 4.1 and 4.2 in [7]. Following Lemma 4.1 in [7], we want to compute $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ such that
@@
Noticing that $\left(\mathbb{L}_{n_{2}}-\mu I\right)^{-1}$ is the convolution operator associated to the function $f_{n_{2},\mu}$ defined in Lemma 5.4, we combine Lemma 4.7 and Lemma 5.4 to obtain that
$2d_{1}(\bar{V},\bar{V}*E)$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2},$
@@
using Lemma 8.1. Then, recalling $C_{n_{2}}(\mu)$ from the proof of Lemma 5.4, one can prove that $\mu\mapsto C_{n_{2}}(\mu)\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$ is increasing in $\mu$ and consequently the maximum is reached at $\mu=\delta_{0}$ when considering $\mu\in\mathcal{J}$. This concludes the inequality for $\mathcal{Z}_{u,3}$. The computations of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ follow directly from the proof of Lemma 4.3. In particular, we also use that $\|(\mathbb{L}-\mu I)^{-1}\|_{2}\leq\frac{1}{1-\frac{c^{4}}{4}-\delta_{0}}$ for all $\mu\in\mathcal{J}$. The rest of the proof follows from [7]. ∎
@@ 5.2 Enclosure of $\theta$
Let $r_{0}>0$ and let $\tilde{u}\in\mathcal{H}$ be a solution to (2) such that $\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}$, where $\bar{u}\in\mathcal{H}$ is constructed as in (20). Choose $\overline{W}\in\pi^{N}\ell^{2}$ such that $\overline{w}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\overline{W})\in\mathcal{H}$. Let $\epsilon_{0},\epsilon>0$ be defined as
@@
where $\overline{w}\in\mathcal{H}$ is an approximation of $\tilde{w}$. In particular, employing Cauchy-Schwarz inequality, we have for the second term
@@
using that $\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}=\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}.$ Applying the Fourier transform, we have for the second norm on the right-hand side that
@@
Furthermore, we have for the third term of $\theta$ that
@@
$\leq\left(\kappa_{1}r_{0}+2c\|\tilde{w}-\overline{w}\|_{2}\right)\left(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\tfrac{r_{0}}{2-c^{2}}\right).$
It remains to estimate $\|\tilde{w}-\overline{w}\|_{2}$. By definition of $\tilde{w}$ in (40), we have
$\|\tilde{w}-\overline{w}\|_{2}=\|-2cD\mathbb{F}(\tilde{u})\partial_{x_{1}}^{2}\tilde{u}-\overline{w}\|_{2}\leq 2c\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\|\partial_{x_{1}}\tilde{u}+\tfrac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
For the second norm on the right-hand side
$\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}\leq\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\left|\partial_{x_{1}}^{2}(\bar{u}-\tilde{u})\right\|_{2}+\frac{1}{2c}\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
For the last term, we estimate
@@
Finally, using $\bar{u}=\gamma^{\dagger}(\bar{U})$ and $\overline{w}=\gamma^{\dagger}(\overline{W})$, we apply Parseval’s identity to get
@@
where $(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}$ for all $(n_{1},n_{2})\in\mathbb{N}_{0}^{2}$. After algebraic manipulation, this concludes the proof. ∎
@@ 6 Results
In this section, we present computer-assisted proofs for the existence of solutions to (2) with Neumann boundary conditions (3). To do so, we begin with approximate solutions obtained by Newton’s method and initialized with carefully constructed guesses based on approximations in the periodic setting described in [32]. To extend the initial guesses in the periodic case to the infinite setting, we project them into the kernel of matrix $\mathcal{T}$, as described in Section 3.1.
@@
Recall that $\nu$ and $N^{\mathrm{FFT}}$ are defined in Section 3.2. We choose $\nu=(1.1,1.1)$ and $N^{\mathrm{FFT}}=(512,512)$. Finitely many cosine coefficients of the approximate solution $\bar{u}$, visualized in Figure 1, can be found in the code in [31]. We report the values for the bounds in floating point numbers displaying finitely many decimals. We have
@@
The stability part follows similarly as the one presented in the proof of Theorem 6.1. In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from one another, and disjoint from the rest of the intervals. ∎
The next theorem and its proof are very similar to the previous ones. However, the stability analysis is more involved, so we choose to present the full proof again.
@@
For the stability part, we obtain $\lambda_{min}\approx-15$, using the notation of Lemma 5.3. Then, we choose $t>\lambda_{min}$ in Lemma 5.5 and compute the bounds of the lemma to compute the Gershgorin intervals $\overline{B_{\epsilon_{j}}(\lambda_{j})}$. We obtained intersecting intervals and could not conclude about the localization of negative eigenvalues. However, we obtained that the union of all the intervals was contained in $(-0.4,\infty)$. This implies that we can improve our choice for $\lambda_{min}$ and choose $\lambda_{min}=-0.45$. Choosing $t=0.5$ in Lemma 5.3 and computing the bounds again, we get tighter radii for the Gershgorin intervals and obtain disjoint intervals on the negative part of $\mathbb{R}$}.
@@
7 Acknowledgements
Matthieu Cadiot was supported by the ANR project CAPPS: ANR-23-CE40-0004-01 and by the FMJH : ANR-22-EXES-0013. Lindsey van der Aalst was partially supported by NWO grant 613.009.132.
@@ 8 Appendix
In the next lemma we provide some technical results about $l_{n}$, where we recall from (6) that
$l_{n_{2}}(\xi_{1})=(2\pi)^{4}(\xi_{1}^{2}+\tilde{n}_{2}^{2})^{2}-c^{2}(2\pi\xi_{1})^{2}+1$
and that $\tilde{n}_{2}=\frac{n_{2}}{2d_{2}}$.
@@
$\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{(2d_{2})^{7}}{6N^{6}}\frac{5}{8(2\pi)^{7}}$
where $N\in\mathbb{N}$ such that $N>\frac{d_{2}c}{\pi}.$